# RiboFlow: Conditional De Novo RNA Sequence-Structure Co-Design via Synergistic Flow Matching

### v2

@@ Appendix D Implementation Details
(a) Distribution of RNA Length
(b) Distribution of Ligand Type
Figure S5: Statistics of the originally collected RiboBind dataset. (a) The number of RNAs within different length ranges, and the vast majority of RNAs exceed the 200-nucleotide limit; (b) The proportion of the top ten categories of ligands.
(a) Distribution of RNA Length
(b) Distribution of Ligand Type
Figure S6: Statistical analysis of the RiboBind dataset, amplified using a dynamic cropping strategy after removing sequences shorter than 30 nucleotides, reveals the following: (a) The amplified dataset exhibits a more balanced distribution of RNA lengths compared to the original RiboBind dataset, making it better suited for model training. (b) The top ten ligand categories in the amplified dataset are still predominantly composed of the seven major ligands (FME, GTP, SAM, PHE, SPM, MPD, and SPD) found in the original dataset. This suggests that the ligand bias introduced by the amplification process remains within an acceptable range.
Figure S7: The results of vina and posescore for the 24 RNA-ligand pairs in the standard evaluation set. For vina, we additionally provide the experimental values under real conditions, indicated by a blue dashed line.
Figure S8: The results of vina and posescore for the 15 RNA-ligand pairs in the few-shot set. For vina, we additionally provide the experimental values under real conditions, indicated by a blue dashed line.

### v3

@@ 9.1 Model Training Details
9.2 Model Inference
We provide the pseudocode in Algorithm 1 for RiboFlow inference to intuitively demonstrate the sampling process of the model.
Algorithm 1 RiboFlow: Inference
1:Input: Ligand structure $\mathcal{M}$, Sampling steps $T$, Initial RNA structure $\mathcal{T}_{0}$ with length NN, Trained model $v_{\theta}$
2:Initialize: $\textit{steps}\leftarrow 0$, $\textit{t}\leftarrow 0$, $\Delta t\leftarrow 1/T$
3:Initialize: complex $\mathcal{C}_{0}$ with RNA $\mathcal{T}_{0}$ and ligand $\mathcal{M}$
4:while $\textit{steps}<T$ do
5:  $x_{t+\Delta t}^{(i)}\leftarrow x_{t}^{(i)}+\bm{v}_{\theta}(x_{t}^{(i)},t;\mathcal{C}_{t})\Delta t$
6:  $r_{t+\Delta t}^{(i)}\leftarrow r_{t}^{(i)}\exp\left(\bm{v}_{\theta}(r_{t}^{(i)},t;\mathcal{C}_{t})\Delta t\right)$
7:  $s_{t+\Delta t}^{(i)}\leftarrow\operatorname{norm}\big(s_{t}^{(i)}+\bm{v}_{\theta}(s_{t}^{(i)},t;\mathcal{C}_{t})\Delta t\big)$
8:  Sample nucleotide type: $c_{t+\Delta t}^{(i)}\sim s_{t+\Delta t}^{(i)}$
9:  $t\leftarrow t+\Delta t$, $\textit{steps}\leftarrow\textit{steps}+1$
10:end while
11:Calculate: $\hat{\Phi}_{1}^{(i)}\leftarrow\bm{v}_{\phi}(\hat{x}_{1}^{(i)},\hat{r}_{1}^{(i)})$
12:Return: Final complex $\mathcal{C}_{1}$
10 Flow Matching
Flow Matching (FM) [28] is a method designed to reduce the simulation required for learning Continuous Normalizing Flows (CNFs), which is a class of deep generative models that generate data by integrating an ordinary differential equation (ODE) over a learned vector field. In this section, we will provide a concise overview of the flow matching approach.
$\frac{d}{dt}\phi_{t}(x)=v_{t}(\phi_{t}(x)),\quad\phi_{0}(x)=x.$ (18)
Once trained, samples can be generated by simulating Equation˜18 using the learned vector field $v_{t}$.
Figure 12: The results of vina score for the 24 RNA-ligand pairs in the sequence-based evaluation set. We additionally provide the experimental values under real conditions, indicated by a blue dashed line.
Figure 13: The results of vina for the 15 RNA-ligand pairs in the few-shot set. We additionally provide the experimental values under real conditions, indicated by a blue dashed line.
11 Limitations and Broader Impact
RiboFlow has certain limitations, which we briefly discuss here. The model’s pre-training heavily depends on RNASolo, a dataset of RNA structures predominantly spanning lengths of 30–200 nucleotides. Consequently, RiboFlow’s generative performance on longer RNA sequences remains constrained. We anticipate, however, that with continued advances in biotechnology, more high-quality RNA crystal structures will become available, which will help alleviate this limitation. In addition, our current evaluations of RNA-ligand binding affinities are entirely computational and have yet to be validated experimentally. To address this, we are actively collaborating with experimental partners to synthesize the designed RNAs and assess their binding affinities through wet-lab experiments. We are optimistic that these future experimental results will provide stronger empirical support for our approach.
As for broader impacts, our work has many potential application areas. For example, designing fluorescent aptamers that bind to specific molecules as biosensors to detect contamination, or targeting specific biological metabolites for disease treatment. Under the premise of sound legal regulation, we believe that RiboFlow will play a significant role in RNA engineering in the future to benefit all of humanity.